# Sum Rate Maximization for NOMA-Assisted Uplink Pinching-Antenna Systems

### v1

@@ VI Conclusion
In this paper, we studied the problem of sum rate maximization in an uplink communication system employing a pinching-antenna. For the NOMA-based transmission scenario, we demonstrated that the associated optimization problem is non-convex and proposed a PSO approach to efficiently obtain a near-optimal solution. The performance of the proposed NOMA scheme was evaluated against two TDMA-based benchmarks: one based on the pinching antenna individually activated for each user, and the other based on the single-pinching-antenna configuration serving all users. Numerical simulations revealed that systems employing a pinching antenna significantly outperform conventional fixed-antenna architectures in terms of achievable sum rate. Additionally, the NOMA-based approach provides superior performance compared to TDMA with a common pinching-antenna configuration, but is outperformed by TDMA when the antenna location is adaptively optimized for each user. These findings highlight the potential of pinching-antenna systems and offer insights into their integration with multiple access techniques for enhanced spectral efficiency.
In this paper, the scenario with a single activated pinching antenna was considered. An important direction for future research is to study the performance of NOMA assisted pinching-antenna systems with multiple activated pinching antennas, where different conclusions about the comparison between NOMA and TDMA schemes might be drawn.

### v2

@@ VI Conclusion
In this paper, we studied sum rate maximization in an uplink NOMA system with a pinching antenna. For the case without QoS constraints, we showed the antenna optimization problem is non-convex and developed a PSO-based method to obtain near-optimal solutions. With QoS guarantees, we proposed an alternating optimization framework combining low-complexity power allocation and a modified PSO algorithm for antenna positioning. Numerical results demonstrated that pinching antennas significantly outperform fixed architectures, and that NOMA consistently outperforms TDMA. These results highlight the promise of pinching-antenna systems for enhancing spectral efficiency, with future work directed toward extending the analysis to systems with multiple pinching antennas and/or multiple waveguides.